OpenStudy (anonymous):

A peanut merchant sells two kinds of peanuts, one for 60 cents per lb and one for 50 cents per lb. How many of each should he mix to get 50lbs worth 52 cents per lb ?

OpenStudy (anonymous):

You are buying 50 pounds of peanuts so x = number of pounds of one kind y = number of pounds of other kind so x + y = 50 Then price wise, you pay .60 per pounds of one kind .50 per pound of other kind and you pay .52 for each of the 50 pounds so .60x + .50y = .52(50) Solve this system x + y = 50 .60x + .50y = .52(50) x + y = 50 60x + 50y = 2600 (multiplied by 100) -60x - 60y = -3000 (multiplied by -60) 60x + 50y = 2600 -10y = -400 added then up y = 40 divided by -10 substitute back in x + y = 50 x + 40 = 50 x = 10 10 pounds of the 60 cent peanuts 40 pounds of the 50 cent peanuts

OpenStudy (anonymous):

thanks second one you've helped me with :)